# Finite abelian groups acting on rationally connected threefolds II: groups of K3 type

### v1

@@ 1. Introduction
The following theorem is the main result of [Lo22] (see Section 2.6 for the definitions of a $G\mathbb{Q}$-Mori fiber space and a $G\mathbb{Q}$-Fano variety).
Theorem 1.3 ([Lo22, Theorem 1.7]).
@@
In [Lo22, Corollary 1.10] it is proven that there are only finitely many isomorphism classes of finite abelian groups of K3 type which faithfully act on a rationally connected threefold. This result follows from two boundedness results. First, there only finitely many isomorphism classes of finite groups that can faithfully act on a K3 surface, see [BH23] for a complete classification. This bounds the group $H$ in the exact sequence (1.2). The second result, which is needed to bound $m$ in (1.2), is the boundedness of the indices of Fano threefolds with canonical singularities. Of course, it follows from the boundedness of Fano threefolds with canonical singularities. Recently, an effective bound equal to $66$ was obtained in [JL25]. Hence the main problem in dealing with groups of K3 type that act on rationally connected threefolds is to bound the number $m$ in (1.2) for each group $H$ that can faithfully act on a K3 surface. It turns out that this number indeed can be effectively bounded.
@@
(5) $\mathbb{Z}/2n\times(\mathbb{Z}/2)^{4}$ $n\geq 1$
@@ Acknowledgements
The work of the first author was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265), supported by the HSE University Basic Research Program, and the Simons Foundation. The work is supported by the state assignment of MIPT (project FSMG-2023-0013). The author is a Young Russian Mathematics award winner and would like to thank its sponsors and jury. The second author achieved his work at the University of Edinburgh, and was supported by the Leverhulme Trust grant RPG-2021-229. The authors thank Ivan Cheltsov, Dmitri Orlov, and Yuri Prokhorov for helpful discussions, and Alexander Kuznetsov for reading the draft of the paper and many useful remarks.
@@ 3. Group action on K3 surfaces
In the exact sequence (3.1), the group $H_{s}$ is isomorphic to one of the following groups:
@@
Let $H$ be a finite abelian group that faithfully acts on a smooth K3 surface. Assume that $H$ is not purely non-symplectic. If $H$ is assumed to be maximal, then $H$ is isomorphic to one of the following groups:
@@
We start with the case of smooth K3 surfaces which are Fermat complete intersections on which the following group $G$ acts faithfully (cf. Lemma 2.17).
@@
Let $C_{1}$ be an elliptic curve with a faithful action of $H^{\prime}=(\mathbb{Z}/2)^{3}$, and $C_{2}$ be an elliptic curve with a faithful action of $H^{\prime\prime}=\mathbb{Z}/2\times\mathbb{Z}/4$ (so that the $j$-invariant of $C_{2}$ is equal to $1728$). Note that $H^{\prime}$ acts on the set of $2$-torsion points $\{a_{1},a_{2},a_{3},a_{4}\}$ of $C_{1}$ transitively, while $H^{\prime\prime}$ acts on the set of $2$-torsion points $\{b_{1},b_{2},b_{3},b_{4}\}$ of $C_{2}$ with two orbits of cardinality $2$, say, it interchanges $b_{1}$ with $b_{3}$ and $b_{2}$ with $b_{4}$. Consider an abelian surface $A=C_{1}\times C_{2}$, which admits the action of $H^{\prime}\times H^{\prime\prime}=(\mathbb{Z}/2)^{4}\times\mathbb{Z}/4$. It follows that $H^{\prime}\times H^{\prime\prime}$ has $2$ orbits of cardinality $8$. Consider the quotient $S=A/\sigma$ where $\sigma=(\sigma_{1},\sigma_{2})$, and $\sigma_{i}$ is the multiplication by 1-1 on $E_{i}$. Then $S$ is a K3 surface with $16$ singular points of type $A_{1}$. From the construction it follows that $S$ admits a faithful action of $H=(H^{\prime}\times H^{\prime\prime})/(\mathbb{Z}/2)=(\mathbb{Z}/2)^{3}\times\mathbb{Z}/4$. One checks that if $C_{1}$ and $C_{2}$ are isogenous then $\rho(S^{\prime})=20$ where $S^{\prime}$ is the minimal resolution of $S$.
@@ 5.4. Group action on a terminal singularity
Assume that
  • either $r>2$ ,
  • or $G_{x}$ is a $2$-group.
Then the lifting $\widetilde{G}_{x}$ is abelian.
Proof.
The first claim is [Lo24, Proposition 7.13]. The second claim follows from the proof of [Lo24, Proposition 7.18]. ∎
Proposition 5.15.
@@
Assume that $x\in X$ is a threefold cyclic quotient singularity of index $r\geq 1$. Assume that $\widetilde{x}\in\widetilde{S}$ is smooth, so $(\widetilde{x}\in\widetilde{S})\simeq(0\in\mathbb{C}^{2})$. Then there exists an injective homomorphism $\widetilde{H}_{x}\to\widetilde{G}_{x}$ which is inverse to the map $\widetilde{G}_{x}\to\widetilde{H}_{x}$ in (5.13). In particular, we have $\widetilde{G}_{x}=\widetilde{H}_{x}\times\mathbb{Z}/r$.
@@ 6. Case $h^{0}(-K_{X})\geq 2$
where $C_{p_{i}}$ and $G_{p_{i}}$ are $p_{i}$-Sylow subgroups of $C$ and $H$, respectively. We will proceed as follows.
@@ 7. Orbits of non-Gorenstein points
If $H=(\mathbb{Z}/2)^{5}$ and $G$ is not of product type, then the non-Gorenstein locus of $X$ must be composed of orbits whose lengths are multiples of $8$. In particular, we have that $k_{i}$ is divisible by $8$ for all $i\in\{1,\dots,l\}$.
@@
If $H=\mathbb{Z}/4\times(\mathbb{Z}/2)^{3}$ and $G$ is not of product type, then the non-Gorenstein locus of $X$ must be composed of orbits whose lengths are multiples of $4$. In particular, we have that $k_{i}$ is divisible by $4$ for all $i\in\{1,\dots,l\}$.
@@ 8. Case $h^{0}(-K_{X})=1$ with half-points
One checks that either the exact sequence (7.2) splits, or $G$ is of product type. Hence we may assume that $G=H\times\mathbb{Z}/m$. It follows that if $G_{2}$ is of product type then $G$ is of product type as well since $\mathfrak{r}(G_{p})\leq 1$ for $p\neq 2$. Hence we may assume that $G=G_{2}$ is a $2$-group.
By Corollary 5.16 and using its notation, we may assume that $\widetilde{x}\in\widetilde{S}$ is smooth. Hence, in the notation of diagram (5.13), we see that $\mathfrak{r}(\widetilde{H}_{x})\leq 2$ and so $\mathfrak{r}(H_{x})\leq 2$. Let $\Sigma=\{x=x_{1},x_{2},x_{3},x_{4}\}$ be the $G$-orbit of $x$ (or, equivalently, its $H$-orbit). Consider the following exact sequence
$0\to H_{x}\to H\to H_{\Sigma}\to 0,$ (8.8)
where $H_{\Sigma}$ is the image of $H$ in the group of permutations of $\Sigma$. Since the action of $H$ on $\Sigma$ is transitive, we have that either $H_{\Sigma}=\mathbb{Z}/4$, or $H_{\Sigma}=(\mathbb{Z}/2)^{2}$. The first possibility $H_{\Sigma}=\mathbb{Z}/4$, $H_{x}=(\mathbb{Z}/2)^{3}$ is not realized as $\mathfrak{r}(H_{x})\leq 2$. Hence we have $H_{\Sigma}=(\mathbb{Z}/2)^{2}$, and either $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$, or $H_{x}=(\mathbb{Z}/2)^{3}$. Similarly, the second subcase is not realized since $\mathfrak{r}(H_{x})\leq 2$. We obtain $H_{\Sigma}=(\mathbb{Z}/2)^{2}$, $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$. In particular, the exact sequence (8.8) splits: $H=H_{x}\times H_{\Sigma}$.
@@
$0\to\mathbb{Z}/2\to\widetilde{H}_{x}\to H_{x}\to 0.$
We claim that $\widetilde{H}_{x}$ is abelian. Indeed, by Proposition 5.14 we see that $\widetilde{G}_{x}$ is abelian (recall that we assumed that $G$ is a $2$-group). Since $\mathfrak{r}(\widetilde{H}_{x})=2$, $\widetilde{H}_{x}$ is abelian and $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$, we have that either $\widetilde{H}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2$, or $\widetilde{H}_{x}=(\mathbb{Z}/4)^{2}$.
We show that if the first possibility is realized then $G$ is of product type. Let $\widetilde{H}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2$. Then $\widetilde{G}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2\times\mathbb{Z}/n$ for some $n\geq 1$, according to Remark 5.17. However, in this case $\mathfrak{r}(G_{x})\leq 2$, and hence $G$ is of product type.
Thus, we have $\widetilde{H}_{x}=(\mathbb{Z}/4)^{2}$. Then $\widetilde{H}_{\widetilde{E}}=H_{E}=\mathbb{Z}/4$ whose generator acts on $\mathbb{C}^{2}$ via the matrix $\begin{pmatrix}\sqrt{-1}&0\\ 0&-\sqrt{-1}\end{pmatrix}$, $\widetilde{H}_{N}=\mathbb{Z}/4$, whose generator acts on $\mathbb{C}^{2}$ via the matrix $\begin{pmatrix}\sqrt{-1}&0\\ 0&\sqrt{-1}\end{pmatrix}$. We have $H_{N}=\mathbb{Z}/2$. Hence, the exact sequence (8.10) splits. As shown above, the exact sequence (8.8) splits as well. Thus, we have
$H=H_{x}\times H_{\Sigma}=H_{N}\times H_{E}\times H_{\Sigma}=\mathbb{Z}/2\times\mathbb{Z}/4\times(\mathbb{Z}/2)^{2}.$
From the local description it follows that $\widetilde{H}_{N}$ preserves the standard form $dx\wedge dy$. It follows that $H_{N}$ acts symplectically on $S$. Hence, there exists a symplectic element of order $4$ in $H=\mathbb{Z}/4\times(\mathbb{Z}/2)^{3}$. However, this contradicts Theorem 3.3. ∎

### v2

@@ 1. Introduction
The following theorem is the main result of [Lo24] (see Section 2.6 for the definitions of a $G\mathbb{Q}$-Mori fiber space and a $G\mathbb{Q}$-Fano variety).
Theorem 1.3 ([Lo24, Theorem 1.7]).
@@
In [Lo24, Corollary 1.10] it is proven that there are only finitely many isomorphism classes of finite abelian groups of K3 type which faithfully act on a rationally connected threefold. This result follows from two boundedness results. First, there only finitely many isomorphism classes of finite groups that can faithfully act on a K3 surface, see [BH23] for a complete classification. This bounds the group $H$ in the exact sequence (1.2). The second result, which is needed to bound $m$ in (1.2), is the boundedness of the indices of Fano threefolds with canonical singularities. Of course, it follows from the boundedness of Fano threefolds with canonical singularities. Recently, an effective bound equal to $66$ was obtained in [JL25]. Hence the main problem in dealing with groups of K3 type that act on rationally connected threefolds is to bound the number $m$ in (1.2) for each group $H$ that can faithfully act on a K3 surface. It turns out that this number indeed can be effectively bounded.
@@
(5) $\mathbb{Z}/2k\times(\mathbb{Z}/2)^{4}$ $k\geq 1$
@@ Acknowledgements
The work of the first author was performed at the Steklov International Mathematical Center and supported by the Ministry of Science and Higher Education of the Russian Federation (agreement no. 075-15-2022-265), supported by the HSE University Basic Research Program, the Simons Foundation, and by the state assignment of MIPT (project FSMG-2023-0013). The first author is a Young Russian Mathematics award winner and would like to thank its sponsors and jury. The second author achieved his work at the University of Edinburgh, and was supported by the Leverhulme Trust grant RPG-2021-229. The authors thank Ivan Cheltsov, Dmitri Orlov, and Yuri Prokhorov for helpful discussions, and Alexander Kuznetsov for reading the draft of the paper and many useful remarks.
@@ 3. Group action on K3 surfaces
In the exact sequence (3.1), if the group $H_{s}$ is abelian, then it is isomorphic to one of the following groups:
@@
Let $H$ be a finite abelian group that faithfully acts on a smooth K3 surface. Assume that $H$ is not purely non-symplectic. If $H$ is assumed to be maximal with these properties, then $H$ is isomorphic to one of the following groups:
@@
We start with the case of smooth K3 surfaces which are Fermat complete intersections on which the following group acts faithfully (cf. Lemma 2.17).
@@
Let $C_{1}$ be an elliptic curve with a faithful action of $H^{\prime}=(\mathbb{Z}/2)^{3}$, and $C_{2}$ be an elliptic curve with a faithful action of $H^{\prime\prime}=\mathbb{Z}/2\times\mathbb{Z}/4$ (so that the $j$-invariant of $C_{2}$ is equal to $1728$). Note that $H^{\prime}$ acts on the set of $2$-torsion points $\{a_{1},a_{2},a_{3},a_{4}\}$ of $C_{1}$ transitively, while $H^{\prime\prime}$ acts on the set of $2$-torsion points $\{b_{1},b_{2},b_{3},b_{4}\}$ of $C_{2}$ with two orbits of cardinality $2$, say, it interchanges $b_{1}$ with $b_{3}$ and $b_{2}$ with $b_{4}$. Consider an abelian surface $A=C_{1}\times C_{2}$, which admits the action of $H^{\prime}\times H^{\prime\prime}=(\mathbb{Z}/2)^{4}\times\mathbb{Z}/4$. It follows that $H^{\prime}\times H^{\prime\prime}$ has $2$ orbits of cardinality $8$. Consider the quotient $S=A/\sigma$ where $\sigma=(\sigma_{1},\sigma_{2})$, and $\sigma_{i}$ is the multiplication by 1-1 on $C_{i}$. Then $S$ is a K3 surface with $16$ singular points of type $A_{1}$. From the construction it follows that $S$ admits a faithful action of $H=(H^{\prime}\times H^{\prime\prime})/(\mathbb{Z}/2)=(\mathbb{Z}/2)^{3}\times\mathbb{Z}/4$. One checks that if $C_{1}$ and $C_{2}$ are isogenous then $\rho(S^{\prime})=20$ where $S^{\prime}$ is the minimal resolution of $S$.
@@ 5.4. Group action on a terminal singularity
Proposition 5.16.
Assume that
  • either $r>2$ ,
  • or $G_{x}$ is a $2$-group.
Then the lifting $\widetilde{G}_{x}$ is abelian.
Proof.
If the first assumption of the proposition holds, then the claim follows by [Lo24, Proposition 7.13]. If the first assumption of the proposition holds, then the claim follows by [Lo24, Proposition 7.18]. ∎
@@
Assume that $x\in X$ is a threefold cyclic quotient singularity of index $r\geq 1$. Assume that $\widetilde{x}\in\widetilde{S}$ is smooth, so $(\widetilde{x}\in\widetilde{S})\simeq(0\in\mathbb{C}^{2})$. Considering $\mathbb{C}^{2}$ as a linear subspace in $\mathbb{C}^{3}$, we see that there exists an injective homomorphism $\widetilde{H}_{x}\to\widetilde{G}_{x}$ which is inverse to the map $\widetilde{G}_{x}\to\widetilde{H}_{x}$ in (5.13). In particular, we have $\widetilde{G}_{x}=\widetilde{H}_{x}\times\mathbb{Z}/m$.
@@ 6. Case $h^{0}(-K_{X})\geq 2$
where $C_{p_{i}}$ and $H_{p_{i}}$ are $p_{i}$-Sylow subgroups of $C$ and $H$, respectively. We will proceed as follows.
@@ 7. Orbits of non-Gorenstein points
If $H=(\mathbb{Z}/2)^{5}$ and $G$ is not of product type, then the non-Gorenstein locus of $X$ is composed of orbits whose lengths are multiples of $8$. In particular, we have that $k_{i}$ is divisible by $8$ for all $i\in\{1,\dots,l\}$.
@@
If $H=\mathbb{Z}/4\times(\mathbb{Z}/2)^{3}$ and $G$ is not of product type, then the non-Gorenstein locus of $X$ is composed of orbits whose lengths are multiples of $4$. In particular, we have that $k_{i}$ is divisible by $4$ for all $i\in\{1,\dots,l\}$.
@@ 8. Case $h^{0}(-K_{X})=1$ with half-points
By Corollary 5.15 and using its notation, we may assume that $\widetilde{x}\in\widetilde{S}$ is smooth. Let $\Sigma=\{x=x_{1},x_{2},x_{3},x_{4}\}$ be the $G$-orbit of $x$ (or, equivalently, its $H$-orbit). Consider the following exact sequences:
$0\to G_{x}\to G\to G_{\Sigma}\to 0,$ (8.8)
$0\to H_{x}\to H\to H_{\Sigma}\to 0.$ (8.9)
where $G_{\Sigma}$ (respectively, $H_{\Sigma}$) is the image of $G$ (respectively, of $H$) in the group of permutations of $\Sigma$, and $G_{x}$ (respectively, $H_{x}$) are stabilizers of $x$ in $G$ (respectively, $H$). Note that $G_{\Sigma}=H_{\Sigma}$.
We will show that in the assumption of the theorem and assuming that $\widetilde{x}\in\widetilde{S}$ is smooth, the $2$-Sylow subgroup $G_{2}$ cannot faithfully act on $X$. So replacing $G$ by $G_{2}$ we will assume that $G$ is a $2$-group.
By Proposition 5.16 we see that $\widetilde{G}_{x}$ is abelian, and hence $\widetilde{H}_{x}$ is abelian. Since $\widetilde{x}\in\widetilde{S}$ is smooth, in the notation of diagram (5.13), we see that $\mathfrak{r}(\widetilde{H}_{x})\leq 2$ and so $\mathfrak{r}(H_{x})\leq 2$. Since the action of $H$ on $\Sigma$ is transitive, we have that either $H_{\Sigma}=\mathbb{Z}/4$, or $H_{\Sigma}=(\mathbb{Z}/2)^{2}$. The first possibility $H_{\Sigma}=\mathbb{Z}/4$, $H_{x}=(\mathbb{Z}/2)^{3}$ is not realized as $\mathfrak{r}(H_{x})\leq 2$. Hence we have $H_{\Sigma}=(\mathbb{Z}/2)^{2}$, and either $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$, or $H_{x}=(\mathbb{Z}/2)^{3}$. Similarly, the second subcase is not realized since $\mathfrak{r}(H_{x})\leq 2$. We obtain $H_{\Sigma}=(\mathbb{Z}/2)^{2}$, $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$. In particular, the exact sequence (8.9) splits: $H=H_{x}\times H_{\Sigma}$.
@@
$0\to\mathbb{Z}/2\to\widetilde{H}_{x}\to H_{x}\to 0.$ (8.13)
Since $\mathfrak{r}(\widetilde{H}_{x})=2$, $\widetilde{H}_{x}$ is abelian and $H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$, we have that either $\widetilde{H}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2$, or $\widetilde{H}_{x}=(\mathbb{Z}/4)^{2}$.
Assume that $\widetilde{H}_{x}=(\mathbb{Z}/4)^{2}$. Then $\widetilde{H}_{\widetilde{E}}=H_{E}=\mathbb{Z}/4$ whose generator acts on $\mathbb{C}^{2}$ via the matrix $\begin{pmatrix}\sqrt{-1}&0\\ 0&-\sqrt{-1}\end{pmatrix}$, $\widetilde{H}_{N}=\mathbb{Z}/4$, whose generator acts on $\mathbb{C}^{2}$ via the matrix $\begin{pmatrix}\sqrt{-1}&0\\ 0&\sqrt{-1}\end{pmatrix}$. We have $H_{N}=\mathbb{Z}/2$. Hence, the exact sequence (8.11) splits. From the local description it follows that $\widetilde{H}_{E}=H_{E}$ preserves the form $dx\wedge dy$ on $\mathbb{C}^{2}$ which descends to $x\in S$. It follows that $H_{E}$ acts symplectically on $S$. Hence, there exists a symplectic element of order $4$ in $H=\mathbb{Z}/4\times(\mathbb{Z}/2)^{3}$. However, this contradicts Corollary 3.9.
Thus we have $\widetilde{H}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2$. Let $\alpha$ and $\beta$ be the generators of the first and the second factors of $\widetilde{H}_{x}=\mathbb{Z}/8\times\mathbb{Z}/2$. Since $\widetilde{H}_{x}/(\mathbb{Z}/2)=H_{x}=\mathbb{Z}/4\times\mathbb{Z}/2$, and we quotient by the element $\sigma$ which acts on $\mathbb{C}^{2}$ via multiplication by 1-1, it follows that $4\alpha$ is equal to $\sigma$. Hence we may assume that $\beta$ is given by the matrix $\begin{pmatrix}{1}&0\\ 0&{-1}\end{pmatrix}$. However, this is a contradiction, because in this case $\beta$ descends to an element which acts non-symplectically on $S$. Indeed, by Corollary 3.9 the element $\alpha$ descends to a non-symplectic automorphism of order $4$ on $S$, and hence the image of $\alpha$ should act on $S$ symplectically. This completes the proof. ∎